\begin{thebibliography}{99}
\bibitem{Fukuda:1998tw}
  Y.~Fukuda {\it et al.}  [Super-Kamiokande Collaboration],
  %``Measurement of a small atmospheric nu/mu / nu/e ratio,''
  Phys.\ Lett.\ B {\bf 433} (1998) 9
  [arXiv:hep-ex/9803006];
  % Y.~Fukuda {\it et al.}  [Super-Kamiokande Collaboration],
  %``Study of the atmospheric neutrino flux in the multi-GeV energy range,''
  % Phys.\ Lett.\ B {\bf 436} (1998) 33
  {\it ibid} {\bf 436} (1998) 33
  [arXiv:hep-ex/9805006];
  Y.~Fukuda {\it et al.}  [Super-Kamiokande Collaboration],
  %``Evidence for oscillation of atmospheric neutrinos,''
  Phys.\ Rev.\ Lett.\  {\bf 81} (1998) 1562
  [arXiv:hep-ex/9807003];
  % Y.~Fukuda {\it et al.}  [Super-Kamiokande Collaboration],
  %``Measurement of the flux and zenith-angle distribution of upward
  %through-going muons by Super-Kamiokande,''
  % Phys.\ Rev.\ Lett.\  {\bf 82} (1999) 2644
  {\it ibid} {\bf 82} (1999) 2644
  [arXiv:hep-ex/9812014];
  S.~Fukuda {\it et al.}  [Super-Kamiokande Collaboration],
  %``Tau neutrinos favored over sterile neutrinos in atmospheric muon  neutrino
  %oscillations,''
  Phys.\ Rev.\ Lett.\  {\bf 85} (2000) 3999
  [arXiv:hep-ex/0009001];
  Y.~Ashie {\it et al.}  [Super-Kamiokande Collaboration],
  %``Evidence for an oscillatory signature in atmospheric neutrino
  %oscillation,''
  Phys.\ Rev.\ Lett.\  {\bf 93} (2004) 101801
  [arXiv:hep-ex/0404034].
\bibitem{Fukuda:1998fd}
  Y.~Fukuda {\it et al.}  [Super-Kamiokande Collaboration],
  %``Measurements of the solar neutrino flux from Super-Kamiokande's first  300
  %days,''
  Phys.\ Rev.\ Lett.\  {\bf 81} (1998) 1158
  [Erratum-ibid.\  {\bf 81} (1998) 4279]
  [arXiv:hep-ex/9805021];
  % Y.~Fukuda {\it et al.}  [Super-Kamiokande Collaboration],
  %``Constraints on neutrino oscillation parameters from the measurement of
  %day-night solar neutrino fluxes at Super-Kamiokande,''
  % Phys.\ Rev.\ Lett.\  {\bf 82} (1999) 1810
  {\it ibid} {\bf 82} (1999) 1810
  [arXiv:hep-ex/9812009];
  % Y.~Fukuda {\it et al.}  [Super-Kamiokande Collaboration],
  %``Measurement of the solar neutrino energy spectrum using neutrino  electron
  %scattering,''
  % Phys.\ Rev.\ Lett.\  {\bf 82} (1999) 2430
  {\it ibid} {\bf 82} (1999) 2430
  [arXiv:hep-ex/9812011];
  S.~Fukuda {\it et al.}  [Super-Kamiokande Collaboration],
  %``Solar B-8 and he p neutrino measurements from 1258 days of
  %Super-Kamiokande data,''
  Phys.\ Rev.\ Lett.\  {\bf 86} (2001) 5651
  [arXiv:hep-ex/0103032];
  % S.~Fukuda {\it et al.}  [Super-Kamiokande Collaboration],
  %``Constraints on neutrino oscillations using 1258 days of  Super-Kamiokande
  %solar neutrino data,''
  % Phys.\ Rev.\ Lett.\  {\bf 86} (2001) 5656
  {\it ibid} {\bf 86} (2001) 5656
  [arXiv:hep-ex/0103033];
  S.~Fukuda {\it et al.}  [Super-Kamiokande Collaboration],
  %``Determination of solar neutrino oscillation parameters using 1496 days  of
  %Super-Kamiokande-I data,''
  Phys.\ Lett.\ B {\bf 539} (2002) 179
  [arXiv:hep-ex/0205075].
\bibitem{Ahmad:2001an}
  Q.~R.~Ahmad {\it et al.}  [SNO Collaboration],
  %``Measurement of the charged current interactions produced by B-8  solar
  %neutrinos at the Sudbury Neutrino Observatory,''
  Phys.\ Rev.\ Lett.\  {\bf 87} (2001) 071301
  [arXiv:nucl-ex/0106015];
  % Q.~R.~Ahmad {\it et al.}  [SNO Collaboration],
  %``Direct evidence for neutrino flavor transformation from neutral-current
  %interactions in the Sudbury Neutrino Observatory,''
  % Phys.\ Rev.\ Lett.\  {\bf 89} (2002) 011301
  {\it ibid} {\bf 89} (2002) 011301
  [arXiv:nucl-ex/0204008];
  % Q.~R.~Ahmad {\it et al.}  [SNO Collaboration],
  %``Measurement of day and night neutrino energy spectra at SNO and constraints
  %on neutrino mixing parameters,''
  % Phys.\ Rev.\ Lett.\  {\bf 89} (2002) 011302
  {\it ibid} {\bf 89} (2002) 011302
  [arXiv:nucl-ex/0204009];
  S.~N.~Ahmed {\it et al.}  [SNO Collaboration],
  %``Measurement of the total active B-8 solar neutrino flux at the Sudbury
  %Neutrino Observatory with enhanced neutral current sensitivity,''
  Phys.\ Rev.\ Lett.\  {\bf 92} (2004) 181301
  [arXiv:nucl-ex/0309004].
\bibitem{Eguchi:2002dm}
  K.~Eguchi {\it et al.}  [KamLAND Collaboration],
  %``First results from KamLAND: Evidence for reactor anti-neutrino
  %disappearance,''
  Phys.\ Rev.\ Lett.\  {\bf 90} (2003) 021802
  [arXiv:hep-ex/0212021];
  T.~Araki {\it et al.}  [KamLAND Collaboration],
  %``Measurement of neutrino oscillation with KamLAND: Evidence of spectral
  %distortion,''
  Phys.\ Rev.\ Lett.\  {\bf 94} (2005) 081801
  [arXiv:hep-ex/0406035].
\bibitem{Apollonio:1997xe}
  M.~Apollonio {\it et al.}  [CHOOZ Collaboration],
  %``Initial results from the CHOOZ long baseline reactor neutrino  oscillation
  %experiment,''
  Phys.\ Lett.\ B {\bf 420} (1998) 397
  [arXiv:hep-ex/9711002];
  % M.~Apollonio {\it et al.}  [CHOOZ Collaboration],
  %``Limits on neutrino oscillations from the CHOOZ experiment,''
  % Phys.\ Lett.\ B {\bf 466} (1999) 415
  {\it ibid} {\bf 466} (1999) 415
  [arXiv:hep-ex/9907037];
  M.~Apollonio {\it et al.},
  %``Search for neutrino oscillations on a long base-line at the CHOOZ  nuclear
  %power station,''
  Eur.\ Phys.\ J.\ C {\bf 27} (2003) 331
  [arXiv:hep-ex/0301017].
\bibitem{Cleveland:1998nv}
  B.~T.~Cleveland {\it et al.},
  %``Measurement of the solar electron neutrino flux with the Homestake
  %chlorine detector,''
  Astrophys.\ J.\  {\bf 496} (1998) 505.
\bibitem{Hampel:1998xg}
  W.~Hampel {\it et al.}  [GALLEX Collaboration],
  %``GALLEX solar neutrino observations: Results for GALLEX IV,''
  Phys.\ Lett.\ B {\bf 447} (1999) 127.
\bibitem{Ahn:2001cq}
  S.~H.~Ahn {\it et al.}  [K2K Collaboration],
  %``Detection of accelerator produced neutrinos at a distance of 250-km,''
  Phys.\ Lett.\ B {\bf 511} (2001) 178
  [arXiv:hep-ex/0103001];
  M.~H.~Ahn {\it et al.}  [K2K Collaboration],
  %``Indications of neutrino oscillation in a 250-km long-baseline experiment,''
  Phys.\ Rev.\ Lett.\  {\bf 90} (2003) 041801
  [arXiv:hep-ex/0212007].
\bibitem{Altarelli:2004za}
  G.~Altarelli and F.~Feruglio,
  %``Models of neutrino masses and mixings,''
  New J.\ Phys.\  {\bf 6} (2004) 106
  [arXiv:hep-ph/0405048].
\bibitem{Maki:1962mu}
  Z.~Maki, M.~Nakagawa and S.~Sakata,
  %``Remarks On The Unified Model Of Elementary Particles,''
  Prog.\ Theor.\ Phys.\  {\bf 28} (1962) 870.
\bibitem{Maltoni:2004ei}
  M.~Maltoni, T.~Schwetz, M.~A.~Tortola and J.~W.~F.~Valle,
  %``Status of global fits to neutrino oscillations,''
  New J.\ Phys.\  {\bf 6} (2004) 122
  [arXiv:hep-ph/0405172].
\bibitem{Pati:1974yy}
  J.~C.~Pati and A.~Salam,
  %``Lepton Number As The Fourth Color,''
  Phys.\ Rev.\ D {\bf 10} (1974) 275;
  %\bibitem{Georgi:1974sy}
  H.~Georgi and S.~L.~Glashow,
  %``Unity Of All Elementary Particle Forces,''
  Phys.\ Rev.\ Lett.\  {\bf 32} (1974) 438;
  %\bibitem{Fritzsch:1974nn}
  H.~Fritzsch and P.~Minkowski,
  %``Unified Interactions Of Leptons And Hadrons,''
  Annals Phys.\  {\bf 93} (1975) 193;
  %\bibitem{Gursey:1975ki}
  F.~Gursey, P.~Ramond and P.~Sikivie,
  %``A Universal Gauge Theory Model Based On E6,''
  Phys.\ Lett.\ B {\bf 60} (1976) 177;
\bibitem{Cabibbo:1963yz}
  N.~Cabibbo,
  %``Unitary Symmetry And Leptonic Decays,''
  Phys.\ Rev.\ Lett.\  {\bf 10} (1963) 531;
  %\bibitem{Kobayashi:1973fv}
  M.~Kobayashi and T.~Maskawa,
  %``CP Violation In The Renormalizable Theory Of Weak Interaction,''
  Prog.\ Theor.\ Phys.\  {\bf 49} (1973) 652.
\bibitem{Mohapatra:1986uf}
  For example, 
  R.~N.~Mohapatra,
  ``Unification And Supersymmetry. The Frontiers Of Quark - Lepton Physics'',
  New York, USA: Springer (2003);
  %\bibitem{Matsuda:2001bg}
  K.~Matsuda, Y.~Koide, T.~Fukuyama and H.~Nishiura,
  %``How far can the SO(10) two Higgs model describe the observed neutrino
  %masses and mixings?,''
  Phys.\ Rev.\ D {\bf 65} (2002) 033008
  [Erratum-ibid.\ D {\bf 65} (2002) 079904]
  [arXiv:hep-ph/0108202];
  %\bibitem{Fukuyama:2002ch}
  T.~Fukuyama and N.~Okada,
  %``Neutrino oscillation data versus minimal supersymmetric SO(10) model,''
  JHEP {\bf 0211} (2002) 011
  [arXiv:hep-ph/0205066];
  %\bibitem{Bajc:2002iw}
  B.~Bajc, G.~Senjanovic and F.~Vissani,
  %``b - tau unification and large atmospheric mixing: A case for non-canonical
  %see-saw,''
  Phys.\ Rev.\ Lett.\  {\bf 90} (2003) 051802
  [arXiv:hep-ph/0210207];
  %\bibitem{Goh:2003sy}
  H.~S.~Goh, R.~N.~Mohapatra and S.~P.~Ng,
  %``Minimal SUSY SO(10), b tau unification and large neutrino mixings,''
  Phys.\ Lett.\ B {\bf 570} (2003) 215
  [arXiv:hep-ph/0303055];
  %\bibitem{Aulakh:2003kg}
  C.~S.~Aulakh, B.~Bajc, A.~Melfo, G.~Senjanovic and F.~Vissani,
  %``The minimal supersymmetric grand unified theory,''
  Phys.\ Lett.\ B {\bf 588} (2004) 196
  [arXiv:hep-ph/0306242];
  %\bibitem{Bajc:2004xe}
  B.~Bajc, A.~Melfo, G.~Senjanovic and F.~Vissani,
  %``The minimal supersymmetric grand unified theory. I: Symmetry breaking  and
  %the particle spectrum,''
  Phys.\ Rev.\ D {\bf 70} (2004) 035007
  [arXiv:hep-ph/0402122];
  %\bibitem{Dutta:2005ni}
  B.~Dutta, Y.~Mimura and R.~N.~Mohapatra,
  %``Neutrino mixing predictions of a minimal SO(10) model with suppressed
  %proton decay,''
  Phys.\ Rev.\ D {\bf 72} (2005) 075009
  [arXiv:hep-ph/0507319].
\bibitem{Albright:1998vf}
  For example, 
  C.~H.~Albright, K.~S.~Babu and S.~M.~Barr,
  %``A minimality condition and atmospheric neutrino oscillations,''
  Phys.\ Rev.\ Lett.\  {\bf 81} (1998) 1167
  [arXiv:hep-ph/9802314];
  %\bibitem{Babu:1998wi}
  K.~S.~Babu, J.~C.~Pati and F.~Wilczek,
  %``Fermion masses, neutrino oscillations, and proton decay in the light of
  %SuperKamiokande,''
  Nucl.\ Phys.\ B {\bf 566} (2000) 33
  [arXiv:hep-ph/9812538];
  %\bibitem{Babu:2001cv}
  K.~S.~Babu and S.~M.~Barr,
  %``Bimaximal neutrino mixings from lopsided mass matrices,''
  Phys.\ Lett.\ B {\bf 525} (2002) 289
  [arXiv:hep-ph/0111215].
\bibitem{Bando:1999km}
  For example, 
  M.~Bando and T.~Kugo,
  %``Neutrino masses in E(6) unification,''
  Prog.\ Theor.\ Phys.\  {\bf 101} (1999) 1313
  [arXiv:hep-ph/9902204];
  %\bibitem{Bando:2000gs}
  M.~Bando, T.~Kugo and K.~Yoshioka,
  %``Mass matrices in E(6) unification,''
  Prog.\ Theor.\ Phys.\  {\bf 104} (2000) 211
  [arXiv:hep-ph/0003220];
  %\bibitem{Bando:2001bj}
  M.~Bando and N.~Maekawa,
  %``E(6) unification with bi-large neutrino mixing,''
  Prog.\ Theor.\ Phys.\  {\bf 106} (2001) 1255
  [arXiv:hep-ph/0109018];
  %\bibitem{Maekawa:2004qj}
  N.~Maekawa,
  %``E(6) unification, large neutrino mixings, and SUSY flavor problem,''
  Prog.\ Theor.\ Phys.\  {\bf 112} (2004) 639
  [arXiv:hep-ph/0402224].
  In SO(10) models, see, 
  %\bibitem{Nomura:1998gm}
  Y.~Nomura and T.~Yanagida,
  %``Bi-maximal neutrino mixing in SO(10)(GUT),''
  Phys.\ Rev.\ D {\bf 59} (1999) 017303
  [arXiv:hep-ph/9807325].
\bibitem{Seesaw}
  T.~Yanagida, in Proceedings of the "{\it Workshop on the Unified Theory and the Baryon Number in the Universe}", Tsukuba, 1979, edited by O. Sawada and A. Sugamoto, KEK Report No. KEK-79-18, p. 95; 
  S.~L.~Glashow, in {\it Quarks and Leptons, Carg\'ese 1979} (Plenum Press, New York, 1980) edited by M.~Levy {\it et. el,} p. 707;
  M.~Gell-Mann, P.~Ramond and R.~Slansky, in Proceedings of the "{\it Supergravity Stony Brook Workshop}", edited by D.~Z.~Freedman and P.~van.~Nieuwenhuizen (North-Holland, Amsterdam, 1979) p. 315;
  %\bibitem{Mohapatra:1979ia}
  R.~N.~Mohapatra and G.~Senjanovic,
  %``Neutrino Mass And Spontaneous Parity Nonconservation,''
  Phys.\ Rev.\ Lett.\  {\bf 44} (1980) 912.
\bibitem{Petcov:1982ya}
  For example, 
  S.~T.~Petcov,
  %``On Pseudodirac Neutrinos, Neutrino Oscillations And Neutrinoless Double
  %Beta Decay,''
  Phys.\ Lett.\ B {\bf 110} (1982) 245;
  %\bibitem{Barbieri:1998mq}
  R.~Barbieri, L.~J.~Hall, D.~R.~Smith, A.~Strumia and N.~Weiner,
  %``Oscillations of solar and atmospheric neutrinos,''
  JHEP {\bf 9812} (1998) 017
  [arXiv:hep-ph/9807235];
  %\bibitem{Mohapatra:1998bp}
  R.~N.~Mohapatra and S.~Nussinov,
  %``Gauge model for bimaximal neutrino mixing,''
  Phys.\ Lett.\ B {\bf 441} (1998) 299
  [arXiv:hep-ph/9808301];
  %\bibitem{Mohapatra:1998ka}
  R.~N.~Mohapatra and S.~Nussinov,
  %``Bimaximal neutrino mixing and neutrino mass matrix,''
  Phys.\ Rev.\ D {\bf 60} (1999) 013002
  [arXiv:hep-ph/9809415];
  %\bibitem{Mohapatra:1999zr}
  R.~N.~Mohapatra, A.~Perez-Lorenzana and C.~A.~de Sousa Pires,
  %``Type II seesaw and a gauge model for the bimaximal mixing explanation  of
  %neutrino puzzles,''
  Phys.\ Lett.\ B {\bf 474} (2000) 355
  [arXiv:hep-ph/9911395];
  %\bibitem{Grimus:2001ex}
  W.~Grimus and L.~Lavoura,
  %``Softly broken lepton numbers and maximal neutrino mixing,''
  JHEP {\bf 0107} (2001) 045
  [arXiv:hep-ph/0105212];
  %\bibitem{Babu:2002ex}
  K.~S.~Babu and R.~N.~Mohapatra,
  %``Predictive schemes for bimaximal neutrino mixings,''
  Phys.\ Lett.\ B {\bf 532} (2002) 77
  [arXiv:hep-ph/0201176];
  %\bibitem{Goh:2002nk}
  H.~S.~Goh, R.~N.~Mohapatra and S.~P.~Ng,
  %``Testing neutrino mass matrices with approximate L(e) - L(mu) - L(tau)
  %symmetry,''
  Phys.\ Lett.\ B {\bf 542} (2002) 116
  [arXiv:hep-ph/0205131];
  %\bibitem{Kuchimanchi:2002yu}
  R.~Kuchimanchi and R.~N.~Mohapatra,
  %``Bimaximal neutrino mixing from a local SU(2) horizontal symmetry,''
  Phys.\ Rev.\ D {\bf 66} (2002) 051301
  [arXiv:hep-ph/0207110];
  %\bibitem{Petcov:2004rk}
  S.~T.~Petcov and W.~Rodejohann,
  %``Flavor symmetry L(e) - L(mu) - L(tau), atmospheric neutrino mixing and  CP
  %violation in the lepton sector,''
  Phys.\ Rev.\ D {\bf 71} (2005) 073002
  [arXiv:hep-ph/0409135].
\bibitem{Giunti:2002ye}
  C.~Giunti and M.~Tanimoto,
  %``Deviation of neutrino mixing from bi-maximal,''
  Phys.\ Rev.\ D {\bf 66} (2002) 053013
  [arXiv:hep-ph/0207096];
  %\bibitem{Giunti:2002pp}
  %C.~Giunti and M.~Tanimoto,
  %``CP violation in bilarge lepton mixing,''
  %Phys.\ Rev.\ D {\bf 66} (2002) 113006
  {\it ibid} {\bf 66} (2002) 113006
  [arXiv:hep-ph/0209169].
\bibitem{Rodejohann:2003sc}
  W.~Rodejohann,
  %``A parametrization for the neutrino mixing matrix,''
  Phys.\ Rev.\ D {\bf 69} (2004) 033005
  [arXiv:hep-ph/0309249];
  %\bibitem{Frampton:2004ud}
  P.~H.~Frampton, S.~T.~Petcov and W.~Rodejohann,
  %``On deviations from bimaximal neutrino mixing,''
  Nucl.\ Phys.\ B {\bf 687} (2004) 31
  [arXiv:hep-ph/0401206].
\bibitem{Datta:2003qg}
  A.~Datta, F.~S.~Ling and P.~Ramond,
  %``Correlated hierarchy, Dirac masses and large mixing angles,''
  Nucl.\ Phys.\ B {\bf 671} (2003) 383
  [arXiv:hep-ph/0306002];
  %\bibitem{Datta:2005ci}
  A.~Datta, L.~Everett and P.~Ramond,
  %``Cabibbo haze in lepton mixing,''
  Phys.\ Lett.\ B {\bf 620} (2005) 42
  [arXiv:hep-ph/0503222];
  %\bibitem{Everett:2005ku}
  L.~L.~Everett,
  %``Viewing lepton mixing through the Cabibbo haze,''
  arXiv:hep-ph/0510256.
\bibitem{Xing:2005ur}
  Z.~z.~Xing,
  %``Nontrivial correlation between the CKM and MNS matrices,''
  Phys.\ Lett.\ B {\bf 618} (2005) 141
  [arXiv:hep-ph/0503200].
\bibitem{Dutta:2005bb}
  B.~Dutta and Y.~Mimura,
  %``Properties of fermion mixings in intersecting D-brane models,''
  arXiv:hep-ph/0512171.
\bibitem{Wolfenstein:1983yz}
  L.~Wolfenstein,
  %``Parametrization Of The Kobayashi-Maskawa Matrix,''
  Phys.\ Rev.\ Lett.\  {\bf 51} (1983) 1945.
\bibitem{Eidelman:2004wy}
  S.~Eidelman {\it et al.}  [Particle Data Group],
  %``Review of particle physics,''
  Phys.\ Lett.\ B {\bf 592} (2004) 1.
\bibitem{Minakata:2004xt}
  H.~Minakata and A.~Y.~Smirnov,
  %``Neutrino mixing and quark lepton complementarity,''
  Phys.\ Rev.\ D {\bf 70} (2004) 073009
  [arXiv:hep-ph/0405088];
  %\bibitem{Smirnov:2004ju}
  A.~Y.~Smirnov,
  %``Neutrinos: '...Annus mirabilis',''
  arXiv:hep-ph/0402264;
  %\bibitem{Raidal:2004iw}
  M.~Raidal,
  %``Prediction Theta(C) + Theta(sol) = pi/4 from flavor physics: A new
  %evidence for grand unification?,''
  Phys.\ Rev.\ Lett.\  {\bf 93} (2004) 161801
  [arXiv:hep-ph/0404046];
  %\bibitem{Frampton:2004vw}
  P.~H.~Frampton and R.~N.~Mohapatra,
  %``Possible gauge theoretic origin for quark-lepton complementarity,''
  JHEP {\bf 0501} (2005) 025
  [arXiv:hep-ph/0407139];
  %\bibitem{Ferrandis:2004vp}
  J.~Ferrandis and S.~Pakvasa,
  %``QLC relation and neutrino mass hierarchy,''
  Phys.\ Rev.\ D {\bf 71} (2005) 033004
  [arXiv:hep-ph/0412038];
  %\bibitem{Kang:2005as}
  S.~K.~Kang, C.~S.~Kim and J.~Lee,
  %``Quark-lepton complementarity with renormalization effects through
  %threshold corrections,''
  Phys.\ Lett.\ B {\bf 619} (2005) 129
  [arXiv:hep-ph/0501029];
  %\bibitem{Cheung:2005gq}
  K.~Cheung, S.~K.~Kang, C.~S.~Kim and J.~Lee,
  %``Lepton flavor violation as a probe of quark-lepton unification,''
  Phys.\ Rev.\ D {\bf 72} (2005) 036003
  [arXiv:hep-ph/0503122];
  %\bibitem{Antusch:2005ca}
  S.~Antusch, S.~F.~King and R.~N.~Mohapatra,
  %``Quark lepton complementarity in unified theories,''
  Phys.\ Lett.\ B {\bf 618} (2005) 150
  [arXiv:hep-ph/0504007];
  %\bibitem{Minakata:2005rf}
  H.~Minakata,
  %``Quark-lepton complementarity: A review,''
  arXiv:hep-ph/0505262;
  %\bibitem{Ohlsson:2005js}
  T.~Ohlsson,
  %``Bimaximal fermion mixing from the quark and leptonic mixing matrices,''
  Phys.\ Lett.\ B {\bf 622} (2005) 159
  [arXiv:hep-ph/0506094];
  %\bibitem{Jarlskog:2005jn}
  C.~Jarlskog,
  %``Ambiguities pertaining to quark-lepton complementarity,''
  Phys.\ Lett.\ B {\bf 625} (2005) 63
  [arXiv:hep-ph/0507212];
  %\bibitem{Joshipura:2005cm}
  A.~S.~Joshipura and A.~Y.~Smirnov,
  %``Quark-Lepton universality and large leptonic mixing,''
  arXiv:hep-ph/0512024;
  %\bibitem{Smirnov:2005kw}
  A.~Y.~Smirnov,
  %``Neutrino masses and mixing,''
  arXiv:hep-ph/0512303.
\bibitem{Jarlskog:1985ht}
  C.~Jarlskog,
  %``Commutator Of The Quark Mass Matrices In The Standard Electroweak Model And
  %A Measure Of Maximal CP Violation,''
  Phys.\ Rev.\ Lett.\  {\bf 55} (1985) 1039.
\end{thebibliography}

\documentclass[preprint,preprintnumbers,amsmath,amssymb]{revtex4}

\usepackage{bm}

\begin{document}

\preprint{APCTP 2005-10}
\preprint{hep-ph/0512294}

\title{Neutrino mixing and CP violation from\\  
    Dirac-Majorana bimaximal mixture and quark-lepton unification}

\author{Junpei Harada}
 \affiliation{Asia Pacific Center for Theoretical Physics, Pohang 790-784, Korea}
 \email{jharada@apctp.org}


\begin{abstract}
We demonstrate that only two ansatz can produce the features of the neutrino mixing angles. The first ansatz comes from the quark-lepton grand unification; $\nu_{Di} = V_{CKM} \nu_{\alpha}$ is satisfied for left-handed neutrinos, where $\nu_{Di}\equiv (\nu_{D1},\nu_{D2},\nu_{D3})$ are the Dirac mass eigenstates and $\nu_{\alpha}\equiv (\nu_e, \nu_\mu, \nu_\tau)$ are the flavour eigenstates. The second ansatz comes from the assumption; $\nu_{Di} = U_{bimaximal} \nu_{i}$ is satisfied between the Dirac mass eigenstates $\nu_{Di}$ and the light Majorana neutrino mass eigenstates $\nu_{i}\equiv (\nu_1, \nu_2, \nu_3)$, where $U_{bimaximal}$ is the $3 \times 3$ rotation matrix that contains two maximal mixing angles and a zero mixing. By these two ansatz, the Maki-Nakagawa-Sakata lepton flavour mixing matrix is given by $U_{MNS} = V_{CKM}^\dagger U_{bimaximal}$. We find that in this model the novel relation $\theta_{sol} + \theta_{13} = \pi/4$ is satisfied, where $\theta_{sol}$ and $\theta_{13}$ are solar and CHOOZ angle respectively. This "Solar-CHOOZ Complementarity" relation indicates that only if the CHOOZ angle $\theta_{13}$ is sizable, the solar angle $\theta_{sol}$ can deviate from the maximal mixing. Our predictions are $\theta_{sol} = 36^\circ$, $\theta_{13} = 9^\circ$ and $\theta_{atm} = 45^\circ$, which are consistent with experiments. We also infer the CP violation in neutrino oscillations. The leptonic Dirac CP phase $\delta_{MNS}$ is predicted as $\sin \delta_{MNS} \simeq A \lambda^2 \eta$, where $A, \lambda, \eta$ are the CKM parameters in Wolfenstein parametrization. In contrast to the quark CP phase $\delta_{CKM} \simeq {\cal O}(1)$, the leptonic Dirac CP phase is very small, $\delta_{MNS} \simeq 0.8^\circ$. Furthermore, we remark that the ratio of the Jarlskog CP violation factor for quarks and leptons is important, because the large uncertainty on $\eta$ is cancelled out in the ratio, $R_J \equiv J_{CKM}/J_{MNS} \simeq 4\sqrt{2} A \lambda^3 \simeq 5 \times 10^{-2}$. 
\end{abstract}

\pacs{}
\keywords{}
\maketitle

In the past several years, our knowledge about the lepton flavour structure has been drastically improved by the progress of neutrino oscillation experiments~\cite{Fukuda:1998tw,Fukuda:1998fd,Ahmad:2001an,Eguchi:2002dm,Apollonio:1997xe,Cleveland:1998nv,Hampel:1998xg,Ahn:2001cq}. Theoretical lepton flavour models have been extensively studied (for a recent review, see e.g.,~\cite{Altarelli:2004za} and references therein), however, there is still no standard answer.
Therefore, trying to understand the lepton flavour structure is a very important theoretical challenge in particle physics today.

In this Letter we demonstrate that only two ansatz can produce the neutrino mixings. Two ansatz will be discussed soon. We find a novel relation $\theta_{sol} + \theta_{13} = \pi/4$ between the solar mixing angle $\theta_{sol}$ and the CHOOZ angle $\theta_{13}$. We call this relation the "Solar-CHOOZ Complementarity" in this Letter. We also infer the CP violation in neutrino oscillations. We remark that the ratio of the Jarlskog CP $J$ factor for quarks and leptons is important, because it does not depend on the CP violation parameter with large uncertainty.

We write the lepton flavour mixing matrix
\begin{eqnarray}
 \nu_{\alpha} = U_{\alpha i} \nu_i,   \label{eq:MNS}
\end{eqnarray}
where $\nu_\alpha \equiv (\nu_e,\nu_\mu,\nu_\tau)$ are the flavour eigenstates and $\nu_i \equiv (\nu_1, \nu_2, \nu_3)$ are the light Majorana mass eigenstates. The $3 \times 3$ unitary matrix $U_{\alpha i}$ is the Maki-Nakagawa-Sakata (MNS) lepton flavour mixing matrix $U_{MNS}$~\cite{Maki:1962mu}. The current status of the neutrino oscillation parameters is elegantly summarized in~\cite{Maltoni:2004ei}.

The first ansatz in our approach comes from the quark-lepton grand unification~\cite{Pati:1974yy}. We write the first ansatz
\begin{eqnarray}
 \nu_{D i} = V_{CKM} \nu_\alpha, \label{eq:1ansatz}
\end{eqnarray}
where $\nu_{Di} \equiv (\nu_{D1},\nu_{D2},\nu_{D3})$ are Dirac mass eigenstates and $\nu_\alpha$ are the flavour eigenstates. The reason that we expect this ansatz is quite simple. For the quark sector, in the Standard Model $u = V_{CKM} u^\prime$ is satisfied for the left-handed quarks, where $u \equiv (u,c,t)$ are the Dirac mass eigenstates and $u^\prime \equiv (u^\prime, c^\prime, t^\prime)$ are the flavour eigenstates. The $V_{CKM}$ is the Cabibbo-Kobayashi-Maskawa (CKM) quark flavour mixing matrix~\cite{Cabibbo:1963yz}. In the quark-lepton grand unification, the Dirac mass structure can be same for quarks and leptons. Therefore, the same Dirac flavour relation would be satisfied between quarks and leptons, and the first ansatz~(\ref{eq:1ansatz}) can be naturally expected. Because of this ansatz, the neutrino mixing angles and the leptonic Dirac CP phase can be given in terms of the CKM quark mixing parameters.

Before discussing the second ansatz, we give some comments on the first ansatz. If you are experimentalists, you can skip this paragraph and go to the second ansatz. We know that the up quark mass matrix equals to the Dirac neutrino mass matrix in SO(10) or $E_6$ grand unification. And the down quark mass matrix relates to the charged lepton mass matrix in SU(5), SO(10) or $E_6$ grand unification. Therefore, in SO(10) or $E_6$ grand unification, the first ansatz~(\ref{eq:1ansatz}) can be naturally expected. However, in these grand unified models, it is well known that the CKM mixing matrix is predicted as $V_{CKM} = {\bf 1}$, because all the standard model fermions (+ right-handed neutrinos) are unified into the same gauge multiplets. We can overcome this difficulty, and construct the realistic models that predict $V_{CKM} = {\bf 1} + \Delta$ by 1) adding the higher dimensional Higgs multiplets~\cite{Mohapatra:1986uf} and/or 2) considering the nonrenormalizable higher dimensional interactions~\cite{Albright:1998vf} and/or 3) using the twisting mechanism in $E_6$ unified models~\cite{Bando:1999km}.  Such explicit model buildings are important studies, however, in this Letter we take a somewhat different view. Instead of specifying scenarios of the unified models, we only impose the first ansatz~(\ref{eq:1ansatz}) that is motivated by the quark-lepton grand unification.

The second ansatz comes from the assumption. We write the second ansatz
\begin{eqnarray}
 \nu_{D i} = U_{bimaximal} \nu_{i},  \label{eq:2ansatz}
\end{eqnarray}
where $\nu_{Di}$ are the Dirac mass eigenstates and $\nu_i$ are the light Majorana mass eigenstates. The $3 \times 3$ rotation matrix $U_{bimaximal}$ contains two maximal mixing angles (1-2 mixing and 2-3 mixing) and a zero 1-3 mixing. This second ansatz means that the dimension 5 effective interaction $y_{ij}L_i L_j H H/M$ is diagonalized by $U_{bimaximal}$, where $L_{i}$ is the SU(2) lepton doublet in the basis of $\nu = \nu_{D i}$ and $H$ is the Higgs doublet. $y_{ij}$ are Yukawa couplings and $M$ is an energy scale to be understood as the right-handed Majorana neutrino mass scale because of the neutrino seesaw mechanism~\cite{Seesaw}.

Before further discussions, we give the reason why we impose the second ansatz by assumption. First of all, even in any grand unified models (for example, SU(5), SO(10) or $E_6$), in general there is no theoretical relation between the Dirac mass eigenstates $\nu_{Di}$ and the light Majorana mass eigenstates $\nu_i$, if we do not impose hypothetical flavour symmetries or texture. Therefore, in this Letter we impose the simplest assumption~(\ref{eq:2ansatz}) that is motivated by the neutrino oscillation experiments~\cite{Fukuda:1998tw,Fukuda:1998fd,Ahmad:2001an,Eguchi:2002dm,Apollonio:1997xe,Cleveland:1998nv,Hampel:1998xg,Ahn:2001cq}, and demonstrate that the assumption~(\ref{eq:2ansatz}) and the first ansatz~(\ref{eq:1ansatz}) are sufficient to produce the neutrino mixings. Theoretical models with flavour symmetries that produce the bimaximal mixing matrix have been studied~\cite{Petcov:1982ya}, however, in this Letter we do not discuss such symmetries.

From Eq.~(\ref{eq:MNS}) and two ansatz~(\ref{eq:1ansatz}) and (\ref{eq:2ansatz}), we can write the neutrino mixing matrix
\begin{eqnarray}
 U_{MNS} = V_{CKM}^\dagger U_{bimaximal}. \label{eq:MNS2}
\end{eqnarray}
This equation shows how neutrino mixing angles and the leptonic Dirac CP phase can be given in terms of the CKM quark mixing parameters. We do not consider the Majorana CP phases in this Letter, which do not affect neutrino oscillations. In our approach, Eq.~(\ref{eq:MNS2}) comes from only two ansatz~(\ref{eq:1ansatz}) and~(\ref{eq:2ansatz}). We comment that the similar relation was studied from different motivations and by different ways~\cite{Giunti:2002ye,Rodejohann:2003sc,Datta:2003qg,Xing:2005ur,Dutta:2005bb}.

In this Letter we consider the CKM mixing matrix in Wolfenstein parametrization~\cite{Wolfenstein:1983yz}. From Eq.~(\ref{eq:MNS2}), the neutrino mixing angles can be expressed in very simple form
\begin{eqnarray}
 \sin^2 2\theta_{atm}  = 1 + {\cal O}(\lambda^4), \quad
 \sin^2 2\theta_{sol} = 1 - 2\lambda^2  + {\cal O}(\lambda^4), \quad
 \sin^2 \theta_{13} = \frac{\lambda^2}{2} + {\cal O}(\lambda^4),  \label{eq:angles}
\end{eqnarray}
where $\lambda \equiv \sin \theta_C$. We find that Eq.~(\ref{eq:angles}) leads to a novel relation between the solar angle $\theta_{sol}$ and the CHOOZ angle $\theta_{13}$
\begin{eqnarray}
 \theta_{sol} + \theta_{13} = \frac{\pi}{4} + {\cal O}(\lambda^3), \label{eq:sol-13}
\end{eqnarray}
where $\lambda^3$ radians $\simeq 0.6^\circ$ for $\lambda = 0.22$~\cite{Eidelman:2004wy}.
This relation is similar to the Quark-Lepton Complementarity~\cite{Minakata:2004xt}, which refers the relation $\theta_{sol} + \theta_C = \pi/4$ between the solar angle $\theta_{sol}$ and the Cabibbo angle $\theta_C (\simeq 13^\circ$~\cite{Eidelman:2004wy}). In this sense, we have named the relation~(\ref{eq:sol-13}) as the "Solar-CHOOZ Complementarity". However, in contrast to the Quark-Lepton Complementarity~\cite{Minakata:2004xt}, Eq.~(\ref{eq:sol-13}) indicates that only if the CHOOZ angle $\theta_{13}$ is sizable, the solar angle $\theta_{sol}$ can deviate from the maximal mixing $\pi/4$. For $\lambda = 0.22$~\cite{Eidelman:2004wy}, our predictions are 
\begin{eqnarray}
 \theta_{atm} = 45^\circ + {\cal O}(\lambda^4), \quad
 \theta_{sol} = 36^\circ + {\cal O}(\lambda^4), \quad
 \theta_{13} = 9^\circ   + {\cal O}(\lambda^3),
\end{eqnarray}
which are consistent with experiments~\cite{Maltoni:2004ei}, and $\lambda^4 (\lambda^3)$ radians $\simeq 0.1^\circ (0.6^\circ)$ for $\lambda = 0.22$~\cite{Eidelman:2004wy}. Furthermore, Eq.~(\ref{eq:sol-13}) and the experimental upper bound on the solar angle $\theta_{sol}$~\cite{Maltoni:2004ei} give the lower bound on the CHOOZ angle $\theta_{13}$
\begin{eqnarray}
 \sin^2 \theta_{sol} < 0.38 \ (0.41) \quad \longrightarrow \quad 7^\circ \ (5^\circ) < \theta_{13},
\end{eqnarray}
where $3 \sigma \ (4 \sigma)$ for the experimental bound~\cite{Maltoni:2004ei}.

We also infer the CP violation in neutrino oscillations. In our model, from Eq.~(\ref{eq:MNS2}), the standard Jarlskog CP violation $J$ factor~\cite{Jarlskog:1985ht} is given by
\begin{eqnarray}
 J_{MNS} \simeq \frac{1}{4\sqrt{2}} A \lambda^3 \eta,   \label{eq:JMNS1}
\end{eqnarray}
where $A$, $\lambda$ and $\eta$ are CKM quark mixing parameters in Wolfenstein parametrization~\cite{Wolfenstein:1983yz}. We can also write the Jarlskog CP violation $J$ factor only with the quantities in Eq.~(\ref{eq:angles})~\cite{Jarlskog:1985ht}
\begin{eqnarray}
 J_{MNS} = \frac{1}{4}\sqrt{\sin^2 2\theta_{sol}} \sqrt{\sin^2 2\theta_{atm}} \sqrt{\sin^2 \theta_{13}}(1-\sin^2 \theta_{13}) \sin \delta_{MNS},   \label{eq:JMNS2}
\end{eqnarray}
where $\delta_{MNS}$ is the leptonic Dirac CP phase.
Therefore, from Eqs.~(\ref{eq:angles}),~(\ref{eq:JMNS1}) and~(\ref{eq:JMNS2}), the leptonic Dirac CP phase is predicted as 
\begin{eqnarray}
 \sin \delta_{MNS} \simeq A \lambda^2 \eta.  \label{eq:phase}
\end{eqnarray}
Eq.~(\ref{eq:phase}) indicates that in contrast to the quark CP phase $\delta_{CKM} = 60^\circ \pm 14^\circ$~\cite{Eidelman:2004wy}, the leptonic Dirac CP phase is very small,
\begin{eqnarray}
 \delta_{MNS} \simeq 0.8^\circ,
\end{eqnarray}
where we have used $\lambda = 0.22$, $A = 0.85$ and $\overline{\eta} = 0.33$ with $\overline{\eta} \equiv \eta (1-\lambda^2/2)$~\cite{Eidelman:2004wy}.

We finally remark that the ratio of the Jarlskog CP violation $J$ factor for quarks and leptons $R_J \equiv J_{CKM}/J_{MNS}$ is important, because the large uncertainty on $\eta$ is cancelled out in the ratio $R_J$,
\begin{eqnarray}
 R_J \equiv \frac{J_{CKM}}{J_{MNS}} \simeq 4 \sqrt{2} A \lambda^3 \simeq 5 \times 10^{-2}, \label{eq:Ratio}
\end{eqnarray}
where $J_{CKM} \simeq A^2 \lambda^6 \eta$~\cite{Jarlskog:1985ht}. Thus, the ratio $R_J$ does not depend on the CP violation parameter $\eta$ and is determined only by the mixing angle parameters $A$ and $\lambda$ with less uncertainty. Although $\delta_{CKM} \gg \delta_{MNS}$, $R_J \ll 1$ is satisfied due to the smallness of the CKM quark flavour mixing angles.

In conclusion, we have demonstrated that only two ansatz~(\ref{eq:1ansatz}) and~(\ref{eq:2ansatz}) are sufficient to produce the neutrino mixing angles, and presented their analytical expressions~(\ref{eq:angles}). We have also discussed the CP violation in neutrino oscillations. If our predictions of the CP violation~(\ref{eq:Ratio}) and the "Solar-CHOOZ Complementarity" relation~(\ref{eq:sol-13}) will be confirmed by the future experiments, we should take our two ansatz~(\ref{eq:1ansatz}) and~(\ref{eq:2ansatz}) seriously and investigate the theoretical origin of the second ansatz~(\ref{eq:2ansatz}).

\acknowledgements
The author thanks the Asia Pacific Center for Theoretical Physics for financial support.

\end{document}